OpenStudy (anonymous):

I need help with a few questions for a long assignment, any takers? I'll medal and fan those who help c:

OpenStudy (anonymous):

I have some information needed to help solve the 2 I need help with but the information needs to be viewed and corrected.

OpenStudy (anonymous):

@iGreen @TheSmartOne @whpalmer4 @sammixboo @uri @Jhannybean @surjithayer @shifuyanli @ikram002p @JFraser @hhelpplzzzz @CausticSyndicalist @linn99123 @Haseeb96 @JoannaBlackwelder @just_one_last_goodbye @welshfella @Phebe @Jaynator495 @IrishBoy123 @Night-Watcher @rishavraj @Lady.Liv1776 @Ilovecake @kidrah69 @JackofallTradez @Surana @nechirwan @10115658 @Musics_life

OpenStudy (surana):

Uh...okay. I'll try. What the heck. I'll go for it.

OpenStudy (anonymous):

thank you so much ;-;

OpenStudy (surana):

I haven't done anything just yet, but you're welcome.

OpenStudy (anonymous):

Rarely anyone responds to my question, so even responding I'm thankful for c: okay so the 2 questions I need help with are Using the accepted values of the processes you've examined, would your estimation of the enthalpy change for the reaction of solid sodium hydroxide in aqueous hydrochloric acid change from the prediction you made in question one? and Give a detailed explanation, using what you know about bonds and forces of attraction, for the enthalpy changes you observed in parts I and II of this lab.

OpenStudy (surana):

Enthalpy...okay. I think I did that yesterday.

OpenStudy (anonymous):

ugh, good, fresh information alright so part 1 has this table Distilled Water Volume 205.0 mL Mass of NaOH 2.535 g NaOH Initial temperature in calorimeter 24.2 °C Final temperature in calorimeter 27.8 °C

OpenStudy (surana):

My response would likely change.

OpenStudy (anonymous):

so you don't think you'd like to help? Enthalpy is kind of a hassle ._.

OpenStudy (surana):

I'm trying. But I got another thing right now so I'm being a bit slow here.

OpenStudy (surana):

@TheYankee ? Could you try this? I'm a bit tied up at present.

OpenStudy (anonymous):

I don't mind, just as long as I can get some sort of help with this c: your help is appreciated. Alright, so what I did for part 1 is as follows: Write out a balanced "equation" for the process you investigated in Part I, including phase symbols. NaOH + H2O -> NaOH Calculate the number of moles of sodium hydroxide dissolved. Show your work. 2.535 grams NaOH * (1 mole / 40 grams) = 0.06338 NaOH moles Calculate the amount of energy involved in this dissolving process. Show your work. q = m c (T2 - T1) 200 mL H2O * 1 g/1 mL = 200 g H2O c = 4.184 J/gC T2 = 27.8 C T1 = 24.2 C 200 g * 4.18 J/gC * (27.8 - 24.2) = 3010 J = Heat gained by H2O Determine the enthalpy change, per mole of sodium hydroxide dissolved. Show your work. 3010 J / 0.06338 NaOH moles = 4.75 * 10^4 J/moles * 1 kJ / 1000 J = 47.5 kJ/moles NaOH

OpenStudy (surana):

Oh. Uh...Okay. Let me see.

OpenStudy (anonymous):

okay c:

OpenStudy (surana):

About as much as I hate to admit it, I don't know. I tried but couldn't come up with something.

OpenStudy (anonymous):

yeah.. This assignment has been bothering me for over a month. I couldn't find much information to help and see if what I had was correct and if I was, how could this help with the 2 questions I don't have anything for o.e well, for future responders, here's part 2 needed to help with the 2 questions Volume of HCl solution 105 mL HCl Volume of NaOH solution 105 mL NaOH Initial temperature in calorimeter 25.2 °C Final temperature in calorimeter 28.2 °C Write out a balanced equation for the reaction you investigated in Part II, including phase symbols. HCI + NaOH -> NaCl + H2O Determine the enthalpy change of this reaction. The mass of the sodium hydroxide and the mass of the hydrochloric acid should be added together. 105 mL HCI + 105 mL NaOH = 210 mL 210 mL * (1 g/mL) = 210 g 210 g * 4.184 j(28.2 - 25.2) = 2475 2475 J * 1 kJ/100 J = 2.475 kJ Determine the number of moles of NaOH. 105 mL NaOH * 1L NaOH/1000 mL NaOH * 0.105 moles NaOH/1L NaOH = 0.0105 NaOH moles Determine enthalpy per mole of NaOH. Show all of your work. 2.475 kJ/0.0105 NaOH moles = 235.7 kJ/mol

OpenStudy (surana):

I'm going to have to depart from this, I'm sorry I couldn't help. Let me tag some people...@misty1212 , @TheYankee and @Nnesha . Maybe they'll succeed where I failed.

OpenStudy (anonymous):

Thank you for responding though, I appreciate your efforts c:

OpenStudy (surana):

:D Here's a medal for trying.

OpenStudy (anonymous):

well thank you c:

OpenStudy (anonymous):

1. No test taking and 2 no mass tagging

OpenStudy (anonymous):

it's not a test, no worries. The mass tagging is for all those who have a better chance at helping me. That and because no one responds. I've waited days at a time for someone to help and repeated questions. Tagging helps aware people realize I actually need help.

OpenStudy (anonymous):

http://prntscr.com/744ro9

OpenStudy (anonymous):

This is an assignment, not a test. A test is given to show your knowledge of a chapter or module.

OpenStudy (anonymous):

@JFraser could you help me out?

OpenStudy (jfraser):

your data seems to make sense to me, you found the energy of the neutralization reaction

OpenStudy (anonymous):

so everything checks out? could you help me with these 2 questions? Using the accepted values of the processes you've examined, would your estimation of the enthalpy change for the reaction of solid sodium hydroxide in aqueous hydrochloric acid change from the prediction you made in question one? and Give a detailed explanation, using what you know about bonds and forces of attraction, for the enthalpy changes you observed in parts I and II of this lab.